solve using elimination or substitution
5/3 d - 3/5 g = 61
3/5 d + 5/3 g = -91
Having trouble clearing the fractions. I used 15 as the LCD and got
25d -9g =915
9d + 25g = -1365
but then I get weird numbers from there...

Answer by Edwin McCravy(20056)

25d -  9g =   915
 9d + 25g = -1365

Multiply the first equation through by 25
Multiply the second equation through by 9

625d - 225g =  22875
 81d + 225g = -12285

Add them term by term and get

706d        =  10590
          d = 15

Substitute in 

   9d + 25g = -1365
9(15) + 25g = -1365
  135 + 25g = -1365
        25g = -1500
          g = -60

Edwin